# Martingale Characterizations of Non-Homogeneous Counting Processes and Their Fractional Variants

### v1

@@ 4. Martingale characterizations of fractional variants
We know that the NGSP can be expressed as the difference between two independent NGCPs, each of which admits a unique representation as a weighted sum of $k$ independent NPPs by Theorem 3.6. Specifically, the NGSP can be written as $\mathcalboondox{S}(t)=\sum_{j=1}^{k}j\mathcal{S}_{j}(t)$, where $\mathcal{S}_{j}(t)=\mathcal{N}_{1j}(t)-\mathcal{N}_{2j}(t)$, with $\{\mathcal{N}_{1j}(t)\}_{t\geq 0}$ and $\{\mathcal{N}_{2j}(t)\}_{t\geq 0}$ denoting independent NPPs having intensity functions $\lambda_{j}(t)$ and $\mu_{j}(t)$ respectively. The NGFSP $\{\mathcalboondox{S}^{\alpha}(t)\}_{t\geq 0}$ is then obtained by time-changing the NGSP $\{\mathcalboondox{S}(t)\}_{t\geq 0}$ with an independent inverse stable subordinator $\{Y_{\alpha}(t)\}_{t\geq 0}$, that is, $\mathcalboondox{S}^{\alpha}(t)=\mathcalboondox{S}(Y_{\alpha}(t))=\sum_{j=1}^{k}j\mathcal{S}_{j}(Y_{\alpha}(t)=\sum_{j=1}^{k}j\mathcal{S}^{\alpha}_{j}(t)$. Since the difference of two martingales (here independent NGFCPs) for a common filtration is a martingale, it follows from Theorem 4.3 and definition of NGFSP that the process $\{X_{j}(t)=\mathcal{S}^{\alpha}_{j}(t)-\big(\Lambda_{j}(Y_{\alpha}(t))-T_{j}(Y_{\alpha}(t))\big)\}_{t\geq 0}$ is a right-continuous martingale with respect to the filtration $\mathcal{F}_{t}^{j}=\sigma\big(\mathcal{S}^{\alpha}_{j}(s):s\leq t\big)\vee\sigma\big(Y_{\alpha}(s):s\geq 0\big)$. The uniform integrability of $\{X_{j}(\tau)\}$, for stopping times $\tau$ satisfying $Y_{\alpha}(\tau)\leq T$, follows from the corresponding property established in Theorem 4.3 and the closure of uniform integrability under linear combinations.
Conversely, suppose there exist simple locally finite point processes $\{\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$ and an independent inverse stable subordinator $\{Y_{\alpha}(t)\}_{t\geq 0}$ such that each $\{X_{j}(t)=\mathcal{S}^{\alpha}_{j}(t)-(\Lambda_{j}(Y_{\alpha}(t))-T_{j}(Y_{\alpha}(t)))\}_{t\geq 0}$ is a right-continuous $\mathcal{F}_{t}^{j}$-martingale, and $\{X_{j}(\tau)\}$ is uniformly integrable for stopping times with $Y_{\alpha}(\tau)\leq T$. Then by Theorem 4.3, each component process $\{\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$ is a NFSP. Consequently, their weighted sum $\{\mathcalboondox{S}^{\alpha}(t)=\sum_{j=1}^{k}j\,\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$ is a NGFSP. This completes the proof. ∎
From Theorem 4.7 and Proposition 4.1, we can obtain the martingale properties of NGSP and NGFSP respectively (see Theorems 7 and 12 in Tathe2024). Next, consider the non-homogeneous tempered generalized space-time fractional Skellam process (NTGSTFSP) which is obtained by time-changing the NGSP with an independent process $\{D_{\beta,\theta}(Y_{\alpha}(t))\}_{t\geq 0}$. Similarly, the non-homogeneous tempered generalized space-fractional Skellam process (NTGSFSP) and the non-homogeneous mixed fractional Skellam process (NMFSP) are obtained by time-changing the NGSP with an independent tempered stable subordinator and an independent inverse mixed stable subordinator, respectively. We now provide the martingale characterization for a NTGSTFSP.
@@
The process $\{\mathcalboondox{S}^{\alpha}_{\beta,\theta}(t)\}_{t\geq 0}$ is a NTGSTFSP if and only if there exist $k$ simple locally finite point processes $\{S_{j\beta,\theta}^{\alpha}(t)\}_{t\geq 0}$, $1\leq j\leq k$ such that $\mathcalboondox{S}_{j\beta,\theta}^{\alpha}(t)=\sum_{j=1}^{k}jS_{j\beta,\theta}^{\alpha}(t)$ and $X_{j}(t)=S_{j\beta,\theta}^{\alpha}(t)-\left(\Lambda_{j}(D_{\beta,\theta}(Y_{\alpha}(t)))-T_{j}(D_{\beta,\theta}(Y_{\alpha}(t)))\right)$ is a right continuous $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$-martingale, where $\Lambda_{j}(t)=\int_{0}^{t}\lambda_{j}(u)du<\infty$ and $T_{j}(t)=\int_{0}^{t}\mu_{j}(u)du<\infty$ for some rate functions $\lambda:[0,\infty)\to[0,\infty)$ and $\mu:[0,\infty)\to[0,\infty)$, $\{Y_{\alpha}(t)\}_{t\geq 0}$ and $\{D_{\beta,\theta}(t)\}_{t\geq 0}$ denote independent inverse stable and tempered stable subordinators respectively, both independent of each other, and $\mathcal{F}_{t}^{j}=\sigma\left(S_{j\beta,\theta}^{\alpha}(s),s\leq t\right)\vee\sigma\left(D_{\beta,\theta}(Y_{\alpha}(s)),s\geq 0\right)$. Moreover, for any $T>0$, $\{X_{j}(\tau),\,\tau\,\,\text{is stopping time such that}(D_{\beta,\theta}(Y_{\alpha}(\tau)))\leq T\}$ is uniformly integrable.
@@
The proof follows from Theorem 4.4 and the fact that the difference of martingales (here independent NTGSTFCPs) defined on a common filtration is a martingale. ∎
Remark 4.4.
The martingale characterization of a NTGSFSP is obtained by putting $\alpha=1$ in Proposition 4.2. The non-homogeneous mixed fractional Skellam process admits a similar characterization, where the process is defined via an inverse mixed stable subordinator.

### v2

@@ 4. Martingale characterizations of fractional variants
We know that a NGSP can be expressed as the difference between two independent NGCPs, each of which admits a unique representation as a weighted sum of $k$ independent NPPs by Theorem 3.6. Specifically, the NGSP can be written as $\mathcalboondox{S}(t)=\sum_{j=1}^{k}j\mathcal{S}_{j}(t)$, where $\mathcal{S}_{j}(t)=\mathcal{N}_{1j}(t)-\mathcal{N}_{2j}(t)$, with $\{\mathcal{N}_{1j}(t)\}_{t\geq 0}$ and $\{\mathcal{N}_{2j}(t)\}_{t\geq 0}$ denoting independent NPPs having intensity functions $\lambda_{j}(t)$ and $\mu_{j}(t)$ respectively. Then the NGFSP $\{\mathcalboondox{S}^{\alpha}(t)\}_{t\geq 0}$ is given by $\mathcalboondox{S}^{\alpha}(t)=\mathcalboondox{S}(Y_{\alpha}(t))=\sum_{j=1}^{k}j\mathcal{S}_{j}(Y_{\alpha}(t))=\sum_{j=1}^{k}j\mathcal{S}^{\alpha}_{j}(t)$, where $\{\mathcal{S}_{j}^{\alpha}(t)\}_{t\geq 0}$ is a NFSP. Therefore for each $\{\mathcal{S}_{j}^{\alpha}(t)\}_{t\geq 0}$, it follows from Theorem 4.8 that the process $\{X_{j}(t)=\mathcal{S}^{\alpha}_{j}(t)-(\Lambda_{j}(Y_{\alpha}(t))-T_{j}(Y_{\alpha}(t)))\}_{t\geq 0}$ is a right-continuous martingale with respect to the filtration $\mathcal{F}_{t}^{j}=\sigma(\mathcal{S}^{\alpha}_{j}(s),s\leq t)\vee\sigma(Y_{\alpha}(s),s\geq 0)$. The uniform integrability of $\{X_{j}(\tau)\}$, for stopping times $\tau$ satisfying $Y_{\alpha}(\tau)\leq T$, follows from the corresponding property established in Theorem 4.8.
Conversely, suppose there exist simple locally finite point processes $\{\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$ and an independent inverse stable subordinator $\{Y_{\alpha}(t)\}_{t\geq 0}$ such that each $\{X_{j}(t)=\mathcal{S}^{\alpha}_{j}(t)-(\Lambda_{j}(Y_{\alpha}(t))-T_{j}(Y_{\alpha}(t)))\}_{t\geq 0}$ is a right-continuous $\mathcal{F}_{t}^{j}$-martingale. Then by Theorem 4.8, each component process $\{\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$ is a NFSP. Consequently, their weighted sum $\{\mathcalboondox{S}^{\alpha}(t)=\sum_{j=1}^{k}j\,\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$ is a NGFSP. This completes the proof. ∎
If the NGFSP is obtained by taking the difference of two independent NGFCPs, the following result provides its martingale characterization.
@@
The process $\{\mathcalboondox{S}^{\alpha_{1},\alpha_{2}}(t)\}_{t\geq 0}$ is a NGFSP if and only if there exist $k$ simple locally finite point processes $\{\mathcal{S}^{\alpha_{1},\alpha_{2}}_{j}(t)\}_{t\geq 0}$, $1\leq j\leq k$ such that $\mathcalboondox{S}^{\alpha_{1},\alpha_{2}}(t)=\sum_{j=1}^{k}j\mathcal{S}^{\alpha_{1},\alpha_{2}}_{j}(t)$ and $X_{j}(t)=\mathcal{S}^{\alpha_{1},\alpha_{2}}_{j}(t)-(\Lambda_{j}(Y_{\alpha_{1}}(t))-T_{j}(Y_{\alpha_{2}}(t))$, $t\geq 0$ is a right continuous $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$-martingale, where $\mathcal{F}_{t}^{j}=\sigma\left(\mathcal{S}_{j}^{\alpha_{1},\alpha_{2}}(s),s\leq t\right)\vee\sigma\left(Y_{\alpha_{1}}(s),s\geq 0\right)\vee\sigma\left(Y_{\alpha_{2}}(s),s\geq 0\right)$, and $\{Y_{\alpha_{1}}(t)\}_{t\geq 0}$ and $\{Y_{\alpha_{2}}(t)\}_{t\geq 0}$ are two independent inverse stable subordinators independent of $\{S_{j}(t)\}_{t\geq 0}$. Moreover, for any $T>0$, {Xj(τ),τis stopping time such that
max(Yα1(τ),Yα2(τ))
T
[…]
is uniformly integrable.
@@
Let $\{\mathcal{M}_{1}^{\alpha_{1}}(t)\}_{t\geq 0}$ and $\{\mathcal{M}_{2}^{\alpha_{2}}(t)\}_{t\geq 0}$ be two independent NGFCPs. By definition, we have $\mathcalboondox{S}^{\alpha_{1},\alpha_{2}}(t)=\mathcal{M}_{1}(Y_{\alpha_{1}}(t))-\mathcal{M}_{2}(Y_{\alpha_{2}}(t))=\sum_{j=1}^{k}j\left(\mathcal{N}_{1j}(Y_{\alpha_{1}}(t))-\mathcal{N}_{2j}(Y_{\alpha_{2}}(t))\right)=\sum_{j=1}^{k}j(\mathcal{N}_{1j}^{\alpha_{1}}(t)-\mathcal{N}_{2j}^{\alpha_{2}}(t))=\sum_{j=1}^{k}j\mathcal{S}_{j}^{\alpha_{1},\alpha_{2}}(t)$, where $\{\mathcal{S}_{j}^{\alpha_{1},\alpha_{2}}(t)\}_{t\geq 0}$ is a NFSP. The result follows from Theorem 4.7 together with the fact that a weighted sum of martingales (here $\{\mathcal{S}_{j}^{\alpha_{1},\alpha_{2}}(t)\}_{t\geq 0}$) with respect to a common filtration (here $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$) is a martingale. ∎
Next, consider the non-homogeneous tempered generalized space-time fractional Skellam process (NTGSTFSP) and its special case NTGSFSP. If they are defined via time-change, then the martingale characterizations of NTGSTFSP and NTGSFSP follow from Proposition  4.1 by replacing $Y_{\alpha}(t)$ with $D_{\beta,\theta}(Y_{\alpha}(t))$ and $D_{\beta,\theta}(t))$ respectively. However, if they are defined as the difference of two independent counting processes, then the martingale characterizations of NTGSTFSP and NTGSFSP follow from Proposition  4.2 by replacing $Y_{\alpha_{i}}(t)$ with $D_{\beta_{i},\theta_{i}}(Y_{\alpha_{i}}(t))$ and $D_{\beta_{i},\theta_{i}}(t))$ for $i=1,2$ respectively. The non-homogeneous mixed fractional Skellam process (NMFSP) admits an analogous characterization when the process is defined via an inverse mixed stable subordinator. Finally, the martingale properties of NGSP, NGFSP, NMFCP, NMFSP, NTGSFCP, NTGSTFCP, NTGSFSP and NTGSTFSP can be obtained from their respective martingale characterizations similar to Corollary 4.1.